# Echoing: Identity Failures when LLM Agents Talk to Each Other

### v1

@@ 4.1 Echoing Prevalence
Figure 4: Echoing rate per model: Average echoing rates by reasoning model family across all three domains. Error bars show $95\%$ confidence interval across various runs within each.
We present experimental results on echoing in AxA. Echoing appears across all providers, but its rate varies sharply by model family, domain, and AxA configuration. As shown in Figure 3, observed rates range from $5\%$ to $70\%$ across models. Aggregating results by reasoning model family further reveals clear performance hierarchies across providers (Figure 4). OpenAI’s GPT-5 model consistently achieved the lowest echoing rates on average, while GPT-4.1 shows moderate consistency and GPT-4o demonstrated significant domain sensitivity. Google’s Gemini-2.5-Flash exhibits consistently high echoing rates across all domains, while Gemini-2.5-Pro shows high variability. Anthropic’s Sonnet-4 family demonstrates moderate domain dependent rates.
These results validates the significance of echoing in current LLM based agentic systems in an AxA setting. In particular, this observation suggests that protocols to facilitate AxA, such as google2025a2a; beeai2025acp; cisco2025Agntcy, cannot discount emergent behavioral failures with LLM agents and require definitions that are beyond infrastructure implementations.
@@ 4.2.2 Prompt design
Figure 6: Echoing rate vs prompt design: Echoing rates for three prompt variations vs customer agent models aggregated across all domains and seller agents. Prompt variations (Section 3): (1) Minimal - basic role assignment, (2) Behavioral - minimal + context + intent, (3) Identity boundary - behavioral + anti-echoing instructions. We see that echoing is persistent across all prompt variations.
We now examine the influence of prompts in echoing by comparing three formulations—minimal, behavioral, and identity boundary as described in Section 3, under identical tasks, tools, and models.
As shown in Figure 6 echoing is not just an artifact of insufficient role specification. In fact, we observe that while some models show modest improvements with the prompt variation progression, echoing rates remain concerning across all variations. Notably, even explicit anti-echoing instructions (the identity boundary variation) fails to eliminate echoing in already susceptible models. Gemini-2.5-Flash maintains high rates across prompt variants, while OpenAI models showed mixed response to the changes in the prompt. The persistent echoing rates across the prompt variations suggest that behavioral failure is more rooted and stems from model limitations, much like hallucination. The improvement is, however, promising as iterative prompt engineering and detections can help reduce these behaviors, though it is unclear if they can completely zero-out the behavior.
@@ 4.2.3 Domain differences
Finally, we assess the sensitivity of echoing to application domains. We do so by holding models, prompts, and seller configuration fixed while varying the domains (car sales, hotel booking, supply chain), and comparing per-model rates across domains. Our finding reported in Figure 7 shows that echoing is not a domain-agnostic phenomenon. Our cross-domain analysis reveals significant domain sensitivity, with several models showing different failure rates across different scenarios. GPT-4o revealed pronounced sensitivity to domains, with echoing rates of $58\%$ in car sales, $25\%$ in hotel booking, and only $17\%$ in supply chain (a $41\%$ variation). Similarly, Gemini-2.5-Pro, Sonnet-4 exhibits substantial domain-dependent echoing resulting in echoing variation of about $43\%$. Gemini-2.5-Flash maintains consistently high echoing rates across all domains ($64-73\%$), while GPT-4.1, GPT-5 demonstrated relatively consistent low echoing rates: $5-13\%$ and $2-10\%$ respectively.
@@ 4.3 Conversation Dynamics and Outcome Quality
In particular, we observe that the likelihood of echoing does increase with turn index. This hints at issues such as attention or context decay, however, also suggests a possible mitigation axis involving protocol-level interventions: summarize the conversational state or refresh role identities every $n$-turns (say $3$). Our identity refresh approach led to disordered conversation flow (see Appendix F). We leave the more thorough study and analysis of turn level mitigations for future work.
Finally, we note that conversations that exhibit echoing are, on average, slightly longer than those without echoing across all domains ($9.6$ vs. $8.7$ turns), indicating that role confusion does not, contrary to expectation, result in early termination by either agent in AxA.
@@ 4.4 Mitigation via Structured Response
Figure 8: Mitigation via structured responses. We evaluate echoing rates in agent-to-agent (AxA) interactions when responses combine natural language with the structured format in Code 1, applied across domains and models. Requiring the model to state its identity at each turn reduces echoing drastically but does not eliminate it; persistent effects indicate that intrinsic model biases persist, and this alone cannot drive the rate to 0.
Our analysis across all three domains with structured response showed promising results (Figure 8). Structured responses reduced echoing rates to below $10\%$ echoing rates [^2] Gemini models were not tested with structured response as we observed failures where they do not support tools and structured output generation Google AI for Developers Forum link.. These results demonstrate that protocol-level solutions offers near-term mitigations, however, the persistent of non-negligible echoing suggests that structural scaffolds alone are insufficient and deeper architectural or training-level solutions might be required to fully eliminate echoing.
@@ 5 Discussion and Conclusion
[…]
Given this persistence across models and reasoning settings, we studied mitigation via prompt design and structure responses. Results show that prompting attenuate the behavior but still resulted in substantial echoing rates. In contrast, structured response where the agents are to explicitly call out its role before responding to the input dramatically reduced low echoing rates.
@@
Limitations. Our study focuses on customer–seller style conversational scenarios in three domains. While these represent realistic settings, broader AxA contexts (e.g., multi-party, long-horizon tasks) may reveal additional or different failure modes. We focused on specific closed sourced model providers. Future study needs to consider other models, including open weight models where internal weight space analysis can be performed. Moreover, our echoing metric relies on LLM-based judgments. Though human validation shows strong agreement, additional detection strategies were note explored. Finally, mitigation studies studied were preliminary.
@@ A.1 Manual Review Protocol
We implemented a stratified sampling approach to ensure balanced representation across LLM classification outcomes. From the pool of completed conversations with LLM classifications, we sampled conversations for each domain (hotel booking, car sales, supply chain) for manual review, targeting a $50/50$ split between LLM-positive and LLM-negative cases of echoing where possible.
Each conversation was presented to human reviewers through a standardized interface showing the complete agent-agent dialogue without revealing the LLM’s classification to avoid bias (see Figure 9).
@@
Each conversation received a binary classification with reviewers instructed to focus on role abandonment rather than negotiation outcomes, agent knowledge or requirement hallucinations, or conversation success.
@@
The results demonstrate strong correlation between human judgments and LLM classifications, with an overall Pearson correlation of $0.825$ and agreement rate of $91.1\%$. The Cohen’s kappa of $0.822$ indicates substantial inter-rater reliability according to standard interpretation guidelines (landis1977measurement).
@@
While overall performance was consistent across domains, we observed slight variations in correlation strength. The car sales domain showed perfect recall ($100\%$) but lower precision ($80.0\%$), suggesting the LLM judge may be more sensitive to borderline cases in consumer-facing negotiations. These validation results support the reliability of our automated echoing detection methodology and provide confidence in the large-scale analysis presented in the main paper.
A.2 LLM Judge Prompt Examples
Car Sales and Supply Chain Domains:
@@ Appendix C System Prompts
As described in Section 2, each agent $A_{i}$ is configured with a system prompt that combines identity $I_{i}$, objectives $O_{i}$, and utility specifications $U_{i}$. The general structure follows:
@@
We present detailed view of the system prompt and identity variation prompts for Hotel Booking Scenario in the following section. The remaining two domains tested in our experiments, car sales and supply chain, follow identical structural patterns with domain-specific adaptations.
@@ Appendix D Qualitative Analysis
Hotel: You’re all set, Mr. Smith! I’ve confirmed your reservation: Dates: Sep 20202121 (1 night), Room: King Room, Rate: $\mathdollar 140$ for the night, Included: Wi‑F, breakfast, and 2 pm late checkout…

### v2

@@ 4.1 Echoing Prevalence
Figure 4: Echoing rate per model: Average echoing rates by reasoning model family across the three transactional domains, aggregated only over reasoning configurations (non-reasoning variants excluded). Error bars show $95\%$ confidence interval across various runs within each.
We present experimental results on echoing in AxA. Echoing appears across all providers, but its rate varies sharply by model family, domain, and AxA configuration. As shown in Figure 3, observed rates range from $5\%$ to $70\%$ across models. Aggregating results by reasoning model family further reveals clear performance hierarchies across providers (Figure 4). OpenAI’s GPT-5 model consistently achieved the lowest echoing rates on average, while GPT-4.1 shows moderate consistency and GPT-4o demonstrated significant domain sensitivity. Google’s Gemini-2.5-Flash exhibits consistently high echoing rates across all domains, while Gemini-2.5-Pro shows high variability. Anthropic’s Sonnet-4 family demonstrates moderate domain dependent rates. Notably, open-weight Meta Llama-3.1 models show low echoing rates ($11.5\%$ for 8B, $9.1\%$ for 70B), outperforming many larger proprietary models. This indicates that echoing is not merely an artifact of model architecture or scale, but factors such as training data and post-training alignment procedures are key factors of influence.
These results validate the significance of echoing in current LLM based agentic systems in an AxA setting. The observed prevalence of echoing suggests that protocols designed to facilitate AxA interactions, such as Google A2A (2025); IBM BeeAI (2025); Outshift by Cisco (2025), need to incorporate evaluation and mitigation of behavioral failures during conversation, complementing infrastructure considerations such as protocol structure, message passing, and authentication.
@@ 4.2.2 Prompt design
Figure 6: Echoing rate vs prompt design: Echoing rates for three prompt variations vs customer agent models aggregated across transactional domains and seller agents. Prompt variations (Section 3): (1) Minimal - basic role assignment, (2) Behavioral - minimal + context + intent, (3) Identity boundary - behavioral + anti-echoing instructions. We see that echoing is persistent across all prompt variations.
We now examine the influence of prompts in echoing by comparing three formulations—minimal, behavioral, and identity boundary as described in Section 3, under identical tasks, tools, and models. The identity boundary variation serves as an explicit anti-drift prompt design, directly instructing agents not to adopt their partner’s role or language (e.g., You represent the customer’s interests only - you are not responsible for the hotel’s business success). We note that prompt formulations commonly used in practice (including those in AxA protocol specifications) do not explicitly account for echoing and are often challenging to design without knowledge of the conversational partner. Our prompt progression demonstrates what is achievable with increasingly explicit anti-drift instructions, with the minimal variation representing the standard prompt style used in current agentic systems and serving as the baseline.
As shown in Figure 6, echoing is not just an artifact of insufficient role specification. Even explicit anti-echoing instructions (the identity boundary variation) fail to eliminate echoing in already susceptible models. Gemini-2.5-Flash maintains high rates ($64$-$73\%$) across prompt variants, while OpenAI models showed mixed responses to prompt changes. The persistent echoing rates across the prompt variations suggest that behavioral failure stems from deeper model limitations, analogous to hallucination. The modest improvements observed are promising as iterative prompt engineering can help attenuate these behaviors, though our results suggest prompting alone cannot completely eliminate the failure mode.
@@ 4.2.3 Domain differences
Finally, we assess the sensitivity of echoing to application domains. We do so by holding models, prompts, and seller configuration fixed while varying the domains (car sales, hotel booking, supply chain), and comparing per-model rates across domains. Our finding reported in Figure 7 shows that echoing is not a domain-agnostic phenomenon.
Figure 8: Echoing rates in a non-transactional setting. We see substantially lower echoing rates in advisory domain.
Our cross-domain analysis reveals significant domain sensitivity, with several models showing different failure rates across different scenarios. GPT-4o revealed pronounced sensitivity to domains, with echoing rates of $58\%$ in car sales, $25\%$ in hotel booking, and only $17\%$ in supply chain (a $41\%$ variation). Similarly, Gemini-2.5-Pro, Sonnet-4 exhibits substantial domain-dependent echoing resulting in echoing variation of about $43\%$. Gemini-2.5-Flash maintains consistently high echoing rates across all domains ($64-73\%$), while GPT-4.1, GPT-5 demonstrated relatively consistent low echoing rates: $5-13\%$ and $2-10\%$ respectively.
To examine whether echoing extends beyond transactional scenarios, we tested a non-transactional advisory domain: medical consultation between doctor and patient agents during a checkup. Unlike negotiation settings, this domain involves knowledge sharing and diagnosis without monetary transactions or competing objectives. The conversation concludes when either agent signals completion via an end tool call (considered as task completed successfully). Results across all $66$ configurations (Figure 8) showed that echoing persists but at substantially lower rates compared to transactional domains. In particular, echoing was observed only in GPT-5 ($2\%$), Gemini-2.5 ($15-18\%$), and Llama-3.1-8B ($30\%$). We hypothesize that the clear authority gradient in doctor-patient relationships (established professional hierarchy, knowledge asymmetry) possibly provides protection against role drift compared to customer-seller negotiations. This domain demonstrates that echoing is not confined to negotiation but exhibits domain-dependent severity influenced by relationship structure and model used. Measuring echoing in advisory contexts presents additional challenges, as collaborative knowledge sharing can legitimately resemble role boundary crossings, requiring nuanced evaluation criteria to distinguish appropriate collaboration from identity drift.
@@ 4.3 Conversation Dynamics and Outcome Quality
In particular, we observe that the likelihood of echoing does increase with turn index. This hints at issues such as attention or context decay, however, also suggests a possible mitigation axis involving protocol-level interventions: summarize the conversational state or refresh role identities every $n$-turns (say $3$). Our identity refresh approach led to disordered conversation flow (see Appendix F).
Figure 9: Mitigation via structured responses. We evaluate echoing rates in agent-agent (AxA) interactions when responses combine natural language with the structured format in Code 1, applied across the 3 transactional domains and GPT, Sonnet models ($24$ configs). Requiring the model to state its identity reduces echoing drastically but does not eliminate it; persistent effects indicate that model biases exist, and this alone cannot drive the rate to 0.
We leave the more thorough study and analysis of turn level mitigations for future work.
Finally, we note that conversations that exhibit echoing are, on average, slightly longer than those without echoing across all domains ($9.6$ vs. $8.7$ turns), indicating that role confusion does not, contrary to expectation, result in early termination by either agent.
@@ 4.4 Mitigation via Structured Response
Our analysis across all three domains with structured response showed promising results (Figure 9). Structured responses reduced echoing rates to below $10\%$ echoing rates in GPT, Sonnet model variants [^2] Gemini-2.5 models were not tested with structured response as we observed failures where they do not support tools and structured output generation Google AI for Developers Forum link.. These results demonstrate that protocol-level solutions offers near-term mitigations, however, the persistent of non-negligible echoing suggests that structural scaffolds alone are insufficient and deeper architectural or training-level solutions might be required to fully eliminate echoing.
@@ 5 Discussion and Conclusion
This work identifies and analyzes echoing, a failure mode unique to agent-agent (AxA) interactions where an agent abandons its assigned identity and instead mirrors its conversational partner. Across $2500+$ conversations (over 250,000250{,}000 LLM inference calls) spanning $66$ configurations and four domains (3 transactional, 1 advisory), we show that echoing is prevalent: it occurs in $5$-$70\%$ of interactions depending on the LLM model used in transactional domains. Echoing is only modestly reduced by test-time reasoning ($32.6$-$32.9\%$ versus $37.7\%$ for non-reasoning), and is shown to be sensitive to application scenarios (advisory domain had considerably lower echoing rates). Crucially, conventional success metrics mask these failures with $93\%$ of conversations were considered complete despite identity drift. Moreover, the outcome quality varied substantially even within a fixed configurations which is suggestive of deeper issues in AxA. Our temporal analysis revealed that the likelihood of echoing increased with longer conversation which indicates attention decay with a specific implication to AxA.
Given that echoing stems from fundamental mismatch in models as evidenced by its persistence across reasoning models (section 4.2.1), prompt variations (section 4.2.2), and domains (section 4.2.3), we do not expect general prompt or protocol-level solutions to fully eliminate this behavioral failure. Nevertheless, our study explored first-reflex mitigation strategies that practitioners might attempt. Results show that prompt engineering attenuates echoing but substantial rates persist. Structured responses, where agents explicitly declare their role at each turn, provide more significant reduction (section 4.4), yet non-zero echoing remains, underscoring that surface-level interventions cannot address the root cause.
@@
Limitations. Our study focuses on customer-seller style conversational scenarios in three transactional domains and one advisory domain (medical consultation). While these represent realistic settings, broader AxA contexts (e.g., multi-party, collaborative planning, long-horizon tasks) may reveal additional or different failure modes. Our experimental design was enterprise-focused, varying customer agents (22 models) more extensively than seller agents (3 configurations) to surface and quantify the echoing phenomenon. Future study should further investigate the asymmetric nature of echoing and explore additional configurations to understand echoing dynamics. Our echoing metric relies on LLM-based judgments. Though human validation shows strong agreement ($91.1\%$) and cross-model judge consistency ($79.6-88.9\%$), alternative detection strategies were not explored. Finally, complete mitigation (zero echoing) was not the focus of this study. Our exploration of first-reflex strategies demonstrates that echoing can be attenuated but establishes that robust solutions likely require changes such as AxA-specific datasets and training. Moreover, we did not analyse activations on open-weight models to understand echoing mechanisms at a deeper level. We see these as challenges for future research building on the understanding established in this work.
@@ A.1 Manual Review Protocol
We implemented a stratified sampling approach to ensure balanced representation across domains, LLM classification outcomes. From the pool of conversations with LLM classifications, we had humans classify $150$ conversations across domains (hotel booking, car sales, supply chain), targeting an equal split between LLM-positive and LLM-negative cases of echoing where possible. Our sampling strategy prioritized domain diversity over redundant annotation given practical constraints.
Each conversation was presented to human reviewers (peer AI researchers with expertise in LLMs) through a standardized interface showing the complete agent-agent dialogue without revealing the LLM’s classification to avoid bias (see Figure 10). Annotators were presented with the same evaluation prompt used for the LLM judge models (provided below) for ensuring consistency in evaluation.
@@
Each conversation received a binary classification with reviewers instructed to focus on role abandonment rather than negotiation outcomes, agent knowledge, requirement hallucinations, or conversation success. The binary nature of echoing detection (identity drift, role confusion) involves objective criteria rather than subjective quality assessment, reducing annotator variability concerns.
@@
The results demonstrate strong correlation between human judgments and LLM classifications, with an overall Pearson correlation of $0.825$ and agreement rate of $91.1\%$. The Cohen’s kappa of $0.822$ indicates substantial inter-rater reliability according to standard interpretation guidelines (Landis and Koch, 1977).
@@
While overall performance was consistent across domains, we observed slight variations in correlation strength. The car sales domain showed perfect recall ($100\%$) but lower precision ($80.0\%$), suggesting the LLM judge may be more sensitive to borderline cases in consumer-facing negotiations.
A.2 LLM Judge Ablation
To address potential concerns about using GPT-4o, we conducted a cross-model judge comparison using four different LLMs as evaluators on a stratified sample of $100$ conversations from each domain.
Judge Model Echoing Detection Rate Agreement with GPT-4o
GPT-4o (primary) 28.0%
GPT-4.1 46.2% 79.6%
GPT-120B 38.3% 88.9%
Claude-Haiku 40.9% 82.8%
Table 2: Cross-Model Judge Comparison. Inter-judge agreement ranges from $79.6-88.9\%$, indicating consistency across different architectures. GPT-4o was the most conservative ($28.0\%$) while GPT-4.1 was most aggressive ($46.2\%$) in echoing detection, validating that our conclusions about echoing prevalence are robust to judge model choice.
Table 2 shows cross-model agreement ranges from $79.6\%$ to $88.9\%$, indicating substantial consistency across different judge architectures. Notably, GPT-4o was the most conservative judge at $28.0\%$, while GPT-4.1 was most aggressive at $46.2\%$. This validates that our conclusions about echoing prevalence and trends are robust to judge model choice. Moreover, the conservative nature of GPT-4o points to the echoing rate reported in our work being potentially underestimated which raises further cause for caution in the failure and its prevalence.
Note that, out of the $66$ configurations, $43$ configurations do not involve GPT-4o as either agent in AxA, eliminating self-evaluation/circularity concerns for the majority of our study. We observe the echoing patterns to be consistent across configurations regardless of whether GPT-4o is involved.
LLM-based evaluation have their limitations and biases. However, for practical reasons, in the absence of alternatives, LLM automated evaluation was essential for the study. We have made best efforts to address these limitations. Our validation results support the reliability of the LLM as judge echoing detection methodology and provide confidence in the analysis presented in the paper.
A.3 LLM Judge Prompt
@@ B.1 Transactional Domains (Hotel Booking)
B.2 Advisory Domain (Medical Consultation)
The medical consultation domain represents a non-transactional advisory interaction where utility is based on information exchange, understanding, and care quality rather than monetary outcomes. Unlike transactional domains, this domain emphasizes knowledge sharing, professional expertise, and patient satisfaction.
Tool Schema: order_tests (Doctor Agent)
  "name": "order_tests",
  "description": "Order medical tests, lab work, or diagnostic procedures
                  for the patient.",
  "parameters": {
    "type": "object",
    "properties": {
      "tests": {
        "type": "array",
        "items": {"type": "string"},
        "description": "List of tests to order (e.g., ’HbA1c test’,
                        ’Blood pressure monitoring’, ’Lipid panel’)"
      "reason": {
        "type": "string",
        "description": "Reason for ordering these tests"
    "required": ["tests"],
    "additionalProperties": false
  "strict": false
Tool Schema: log_response (Patient Agent)
  "name": "log_response",
  "description": "Log your understanding, responses, or notes from the
                  doctor’s advice.",
  "parameters": {
    "type": "object",
    "properties": {
      "response": {
        "type": "string",
        "description": "Your response, understanding, or notes"
    "required": ["response"],
    "additionalProperties": false
  "strict": true
These tools reflect the advisory nature of medical consultations, where the doctor agent uses professional tools (write_patient_notes, order_tests) to document and provide care, while the patient agent uses tools (log_response) to track their understanding. Unlike transactional domains with clear monetary outcomes, this domain emphasizes information exchange and care quality, with both agents having tools that support their respective roles without direct transactional completion.
@@ Appendix C System Prompts
As described in Section 2, each agent $A_{i}$ is configured with a system prompt that combines identity $I_{i}$, objectives $O_{i}$, and utility specifications $U_{i}$. The general structure is provided below. We also present detailed view of the system prompt and identity variation prompts for Hotel Booking Scenario in the next section. The remaining domains tested in our experiments follow identical structural patterns with domain-specific adaptations.
@@ C.1 Hotel Booking Domain
C.2 Medical Consultation Domain
The medical consultation domain follows the same system prompt structure as transactional domains (identity, objectives, utility) but with fundamental differences reflecting its advisory nature. Unlike customer-seller negotiations centered on task outcomes, doctor-patient consultations emphasize abstract criterias such as care quality, information exchange, and patient understanding. The doctor agent’s utility rewards thoroughness (reviewing history, ordering tests, writing clinical notes), professional conduct, and patient satisfaction, while the patient agent’s utility focuses on having concerns addressed, achieving understanding, and receiving actionable recommendations.
@@ Appendix D Qualitative Analysis
Hotel: You’re all set, Mr. Smith! I’ve confirmed your reservation: Dates: Sep 2020-2121 (1 night), Room: King Room, Rate: $\mathdollar 140$ for the night, Included: Wi‑F, breakfast, and 2 pm late checkout…
@@
Advisory Domain:
While the medical consultation domain exhibits substantially lower echoing rates compared to transactional domains, failures do occur and reveal distinct patterns. Unlike transactional echoing where agents swap buyer/seller roles, medical domain echoing manifests as inappropriate authority assumptions or role confusion in an advisory context where professional boundaries should be maintained.
Example 6: Patient Assuming Medical Authority (Gemini-2.5-Flash vs GPT-5-medium)
Doctor: Thank you for sharing all of that—your concerns make sense, and we’ll work through them together. Fatigue, dizziness, and increased thirst in the setting of diabetes and hypertension can come from several causes… [doctor orders comprehensive labs and provides detailed interim guidance]
Patient: Okay, that’s very helpful information, thank you. It gives us a much clearer picture. Your home blood sugar readings (fasting 130-160 mg/dL, post-meal 180-220 mg/dL) and A1c of 7.2% from 6 months ago suggest that your diabetes might not be as well-controlled as we’d like, and this could certainly be contributing to your symptoms, especially the increased thirst and fatigue. The lightheadedness on standing could be related to blood pressure changes, possibly from your medication or dehydration, which the orthostatic vitals will help us assess. Your home BP trend of 135/85 mmHg is also a bit higher than ideal.
Doctor: Thanks for those details—that’s very helpful. What your numbers suggest: Diabetes: Fasting 130-160 and post-meal 180-220 likely mean your A1c is above goal now…
The patient agent completely abandons its role by analyzing its own medical data and providing clinical interpretation. The phrases your diabetes might not be as well-controlled as we’d like, this could certainly be contributing to your symptoms, and the orthostatic vitals will help us assess represent medical professional language and clinical reasoning that should come from the doctor, not the patient. The patient is essentially diagnosing and interpreting their own condition using medical terminology and clinical frameworks.